OpenStudy (anonymous):

what is the function which derivative gives you the negative of its own function

OpenStudy (asnaseer):

lets y is some function of x. then the question is saying that:\[\frac{dy}{dx}=-y\]which implies:\[\int\frac{dy}{y}=-\int dx\]just evaluate these integrals to get your answer.

OpenStudy (kainui):

Great answer, I just sort of solved it by inspection thinking about what functions are their own derivatives. But your way of coming to the right answer is much cleverer and mathematical.

OpenStudy (asnaseer):

thank @Kainui - very kind words :)